OpenStudy (anonymous):

Perform the indicated operation. 2 3/8 · 5 3/4 19/46 13 19/46 13 21/32

OpenStudy (cpinder):

Convert them to fractions, 2 3/8 = 19/8 and 5 3/4 = 23/4. Then do top times top, bottom times bottom. 19*23=437 and 8*4=32. 437/32. Simplify this and you'll get 13 21/32.